# MyFramelinecolor=black, outerlinewidth=2pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt, backgroundcolor=gray!40

### v1

@@ Abstract
Despite being a powerful concept, distributed shared memory (DSM) has not been made practical due to the extensive synchronization needed between servers to implement memory coherence. This paper shows a practical DSM implementation based on the insight that the ownership model embedded in programming languages such as Rust automatically constrains the order of read and write, providing opportunities for significantly simplifying the coherence implementation if the ownership semantics can be exposed to and leveraged by the runtime. This paper discusses the design and implementation of DistR, a Rust-based DSM system that outperforms the two state-of-the-art DSM systems GAM and Grappa by up to 2.64$\times$ and 29.16$\times$ in throughput, and scales much better with the number of servers.
@@ 1 Introduction
When developing a DSM system on top of an ownership-based language like Rust, SWMR semantics are inherently embedded in any Rust program by design. Effortlessly extracting such information becomes possible with basic compiler support, sparing developers from the need for code rewriting. Utilizing the SWMR semantics from the program leads to a considerably simplified process for accessing data in DSM. In the case of a write access, the ownership type ensures exclusive access to the data. Consequently, DistR can move the data to the requesting machine, performing the write there without explicitly invalidating its copies on other machines. In the case of a read access, data can be efficiently replicated to (and cached in) each requesting machine, benefiting from the compiler-provided assurance of freedom from concurrent writes.
This paper presents DistR, an efficient Rust-based DSM implementation that enables object-level concurrent accesses by leveraging the SWMR semantics made explicit by Rust’s ownership type. DistR automatically turns a single-machine Rust program into a DSM-based distributed version without requiring code rewriting. While extracting the ownership semantics appears straightforward, leveraging it to implement a distributed coherence protocol correctly and efficiently presents two main challenges.
@@
To tackle these issues, DistR builds a global heap spanning multiple servers based on the idea of partitioned global address space [21]. Each object in the heap has a unique global address in the address space, which can be used for accessing the object from any server. DistR re-implements Rust’s memory management constructs to allocate objects in the global heap. Given that a server can have cached objects (to accelerate reads), DistR carefully crafts an ownership-based cache coherence protocol upon the global heap abstraction to achieve both memory coherence and efficiency (§4.1.1).
@@
In a nutshell, our coherence protocol leverages the ownership semantics to eliminate the need for explicit cache invalidation. It allows multiple readers to fetch a copy of the object from its host server and cache it, but disallows any change to the global address and the value of the object. When a write access occurs, it must first borrow the ownership, at which point DistR moves the object in the global heap to a new address on the server issuing the write. The address change of the object automatically invalidates cache copies that use the stale address and triggers the subsequent readers to update the cache by fetching the object from its latest address.
@@
The second challenge is how to support transparency in programming. Rust’s standard libraries and programs were originally built for running on a single machine, and they cannot deal with distributed resources in a cluster. For example, a Rust program running on server A cannot spawn a thread on another server B, let alone synchronize threads between A and B. To enable a Rust program to run as is under DistR, we provide distributed threading utilities by restructuring critical elements of the Rust standard library, including threading, communication channels, and shared-state locks (§4.1.2). Our adapted libraries offer the same interfaces, making them compatible with single-machine Rust programs, but internally invoke our distributed scheduler, which determines where to run the thread and facilitates cross-server synchronization. We built them atop the ownership-based memory model, enabling the DistR runtime to safely pass references of objects between threads and automatically fetch the value from the global heap upon dereferencing.
@@
With our programming abstractions, a Rust application can start on a single server and gradually spawn its threads to other servers. Under the hood, DistR employs a runtime to manage distributed physical compute and memory resources for the application. The runtime runs as a process on each node in the cluster, and they work cooperatively for cross-server memory allocation and thread scheduling. The runtime prioritizes the current server for object allocation and thread creation, but it will schedule the resource allocation request to another server under memory pressure (§4.2.1). To make cluster-wise decisions such as deciding the target server for global memory allocation and thread creation, DistR has a global controller that is launched together with the application. The global controller communicates with DistR runtime on each node to collect resource usage information and applies adaptive policies to achieve load balance (§4.2.2).
@@
Results. We evaluated our system on four real-world applications in an eight-node cluster. Our evaluation demonstrated an average of 2.02$\times$ and 9.48$\times$ (up to 2.64$\times$ and 29.16$\times$) speedup compared with two state-of-the-art DSM systems GAM and Grappa, respectively. Furthermore, DistR incurred a mere 2.42% slowdown compared to the original Rust program on a single machine with sufficient resources.
@@ 4 Design
Figure 2: Design overview of DistR.
DistR is an efficient DSM framework atop the Rust programming language. As shown in Figure 2, it consists of Rust-based programming abstractions for DSM (§4.1) and a runtime (§4.2) that manages distributed physical resources.
DistR is compatible with standard Rust. Listing 2 illustrates how the accumulator (shown in Listing 1) runs on DistR distributively without requiring code rewriting. The program starts running on a single machine A and the DistR runtime gradually allocates its memory and spawns new threads on different machines. Specifically, Lines 10LABEL:lst:dist-acc:line:acc-end create Accumulator a and b where a.val and b are in the global heap. We use a global allocator to allocate objects in the global address space and hence these objects may be allocated on a different server. Line 15 synchronously adds b to a by fetching both values a.val and b to A’s local memory (if they are allocated somewhere else). Line LABEL:lst:dist-acc:line:remote-add spawns a new thread and ships the function closure to perform add asynchronously. This thread will be scheduled on a different server B if A’s compute power has been saturated. In this case, DistR performs shallow copying and only ships the pointers stored in a and b to B without actually moving objects in the global heap. The newly-created thread relies on the DistR runtime to detect data locations and fetch objects upon dereferencing.
@@
Listing 2: DistR seamlessly transforms an unmodified accumulator implemented in Rust into a distributed version.
4.1 DistR Programming Abstraction
DistR provides each thread with a local stack and abstracts distributed memory as a shared global heap. Each server allocates thread stacks and backs one partition of the global heap with its physical memory. DistR re-implemented core memory management constructs including Box, &, and &mut for transparent heap access. This approach hides the complex details of memory allocation/deallocation, moving objects, and coherence maintenance (§4.1.1). DistR supports distributed threading and synchronization by adapting Rust’s standard libraries atop the core language constructs (§4.1.2). Furthermore, DistR offers affinity annotations that allow developers to build more efficient applications by expressing data affinity semantics (§4.1.3).
@@ 4.1.1 Memory Management
Figure 3: The address space layout of DistR. The stack is private to each thread but they share an aligned address space to ease migration, while the heap is globally shared and partitioned across servers.
Next, we discuss how DistR (re)implements the memory-related language constructs in Rust to achieve memory safety and memory coherence.
Address Space. As shown in Figure 3, DistR maintains an identical address space layout on all servers. It exposes distributed memory as a coherent shared heap to applications. Embracing the idea of partitioned global address space (PGAS) [21], it partitions the heap space and assigns each server a unique address range. The stack, in contrast, is private to each thread. However, DistR aligns the stack space on each server and pads stacks to avoid overlapping. This streamlines thread migration between servers as it allows a thread to keep its private stack address unchanged when being moved.
Coherence Protocol in a Nutshell. For efficiency, DistR employs a call-by-reference model for newly created threads. Upon creation of a thread, the DistR runtime only passes references or Box pointers to objects to the newly created thread. Upon dereferencing, objects are fetched to the server where the thread is executed.
@@
Note that this is a general protocol that covers the case that the object is on the same server that issues the write—as long as the server moves the object into a different location in the global heap, no other servers can read the stale copies of the object. However, this is not efficient as each local write requires moving the object to a new address. To address this inefficiency, DistR employs a pointer-coloring technique, inspired by the designs of many concurrent garbage collectors [1, 53]. Discussed at the end of this subsection, this technique offers a more efficient solution for handling local writes.
Figure 4: DistR repurposes Rust pointers and references to contain a global heap address and an extension field for its coherence protocol.
Pointer Layout. In order to support this protocol, each pointer must remember not only the object’s global address, but also the address of the cached copy in a server’s local cache (to avoid redundant remote fetches). As such, we modify Rust’s pointer structure, as illustrated in Figure 4. DistR internally extends each Rust Box pointer and reference with an additional 64-bit field, which is used differently for read and write access. At a high level, the field records the address of the cached copy for faster read accesses; for write accesses, this field records the address of the object’s owner for post-write synchronization. Additionally, DistR reserves the highest 16 bits in the global address field as “color” bits. These bits record the version number of the pointer and play a crucial role in DistR’s efficient handling of local writes.
Next, we discuss how DistR reimplements Rust’s ownership operations to realize the distributed coherence protocol. For ease of presentation, this subsection focuses on a simplified version of the protocol. A complete coherence protocol and its proof of memory coherence are available in Appendix B.
@@ 2 Function DerefMut($m$):
Mutable Borrow. Mutable borrow creates a mutable reference that holds exclusive access to the referenced object for writing. Algorithm 1 outlines the procedures for both dereferencing and dropping a mutable reference. When performing dereferencing, DistR first checks the object’s location (Line 1) and performs direct access if the object’s address belongs to the heap partition of the machine A that executes the access. Otherwise, DistR moves it to A’s heap partition (as opposed to caching it) (Line 1). The move, conducted in the following three steps, changes the object’s global address. DistR (1) copies the object into A’s heap at an address $p$, (2) updates the mutable reference with the address $p$, and (3) asynchronously requests the remote server that previously stored the object to deallocate the original object.
A challenge arises with its original owner Box, which now becomes a dangling pointer, pointing to an invalid memory location. Fortunately, the integrity of the system is maintained by the single-writer invariant (referenced as Invariant 3). This invariant ensures that while the mutable reference remains alive, no other entity, including the original owner, can access the data. To ensure correctness, when this new reference is dropped, DistR synchronously updates the original owner Box, redirecting it to the new address $p$ (Line 1). As a result, the original owner always possesses the latest view of the object. Additionally, all modifications made through this mutable reference are visible in all subsequent accesses, as they necessitate borrowing permission from the updated owner Box. The single-writer invariant also eliminates the possibility of simultaneous updates to the owner, ensuring that updating the owner is free from concurrency issues.
@@ 26             Atomic {
Immutable Borrow. Immutable borrowing allows concurrent reads to the same object from immutable references on the same or different servers. As detailed in Algorithm 2, DistR handles the dereferencing of immutable references by first checking the object’s location (Line 2). For remote objects, DistR creates a local copy in the per-node read-only “cache” and records its local address in the reference’s extension field (see Figure 4). This preserves the original global address of the object, ensuring that any new immutable reference—whether it is derived from the owner Box or from another immutable reference—can always access the original object from the global heap.
@@
As opposed to being a separate memory space, our “cache” provides a “virtual” aggregation of all local copies maintained on each server. These copies reside in the regular heap, managed by a per-node hashmap $H$. This hashmap maps each global address to a pair of its local address and the number of local immutable references to the local copy. To prevent redundant copies of an object on the same server, DistR checks the hashmap $H$ before creating a new local copy (Line 2). If a local copy is already present, DistR increments its reference count in $H$ and updates the extension field in the immutable reference to point to this copy (Lines 22). If no existing copy is found, a new one is created (Lines 22). Since the hashmap uses objects’ global addresses as keys, if an object has been modified by another server since its last read, its global address must have changed, making cache lookup fail even if a (stale) local copy exists.
@@
DistR actively updates the reference count of each local copy when an immutable reference is either dereferenced or dropped, as outlined in Lines 2 and 2. Utilizing these counts, the DistR runtime periodically scans the “cache” and lazily reclaims unreferenced copies (i.e., those with a zero reference count) under memory pressure (§4.2.1). This mechanism, in conjunction with the safe borrowing invariant (2), prevents the local cache from memory leaks or illegal accesses.
@@
Owner Access without Borrow. DistR treats a direct memory access via the owner Box as a pair of mutable/immutable borrow and return. Depending on the reference type, DistR uses the extension field of Box accordingly and executes the read/write dereferencing logic. A special case arises when a mutable owner is immutably borrowed and becomes immutable until all borrowed references return. In this case, the owner can only cache the object during the borrow and delay the move until the borrow finishes. This would not create any correctness issues because the owner cannot be used for write access during this period.
@@
Ownership Transfer. Similar to Rust, DistR does not move the actual value during the transfer and only copies the Box pointer. DistR additionally checks and resets the pointer’s extension field and frees the cached copy in the executing machine’s cache to avoid cache leakage.
@@
Memory Deallocation. Like Rust, DistR tracks the lifetime of an object via its owner. Given that ownership transfer is implemented by only evicting the cached copy of the object (without changing its global presence), the memory safety of DistR’s global heap is preserved by the singular owner invariant (1). In other words, DistR still guarantees that when an object’s owner goes out of scope, the object must be unreachable (and dead) and can be safely deallocated.
@@
Consistency Model. Our protocol, together with Rust’s ownership model, offers sequential consistency for cross-server memory accesses in safe Rust programs (i.e., following the original Rust, no guarantees can be provided when Rust Unsafe is used), which is a strong consistency order. Therefore, it allows any safe Rust program to preserve its memory consistency on DSM. Sequential consistency necessitates a coherent memory system, requiring not only the SWMR invariant but also the data-value invariant [57]. In simple terms, the data-value invariant requires that the latest write to a value is immediately visible to subsequent readers. As discussed earlier, DistR’s protocol moves an object upon a write and updates the owner immediately. Therefore, the latest value is globally visible after each mutable borrow finishes. Subsequent read accesses, either in the Owned state or the Shared state, are hence guaranteed to see the moved object and read its latest value.

### v2

@@ Abstract
Despite being a powerful concept, distributed shared memory (DSM) has not been made practical due to the extensive synchronization needed between servers to implement memory coherence. This paper shows a practical DSM implementation based on the insight that the ownership model embedded in programming languages such as Rust automatically constrains the order of read and write, providing opportunities for significantly simplifying the coherence implementation if the ownership semantics can be exposed to and leveraged by the runtime. This paper discusses the design and implementation of DRust, a Rust-based DSM system that outperforms the two state-of-the-art DSM systems GAM and Grappa by up to 2.64$\times$ and 29.16$\times$ in throughput, and scales much better with the number of servers.
@@ 1 Introduction
When developing a DSM system on top of an ownership-based language like Rust, SWMR semantics are inherently embedded in any Rust program by design. Effortlessly extracting such information becomes possible with basic compiler support, sparing developers from the need for code rewriting. Utilizing the SWMR semantics from the program leads to a considerably simplified process for accessing data in DSM. In the case of a write access, the ownership type ensures exclusive access to the data. Consequently, DRust can move the data to the requesting machine, performing the write there without explicitly invalidating its copies on other machines. In the case of a read access, data can be efficiently replicated to (and cached in) each requesting machine, benefiting from the compiler-provided assurance of freedom from concurrent writes.
This paper presents DRust, an efficient Rust-based DSM implementation that enables object-level concurrent accesses by leveraging the SWMR semantics made explicit by Rust’s ownership type. DRust automatically turns a single-machine Rust program into a DSM-based distributed version without requiring code rewriting. While extracting the ownership semantics appears straightforward, leveraging it to implement a distributed coherence protocol correctly and efficiently presents two main challenges.
@@
To tackle these issues, DRust builds a global heap spanning multiple servers based on the idea of partitioned global address space [21]. Each object in the heap has a unique global address in the address space, which can be used for accessing the object from any server. DRust re-implements Rust’s memory management constructs to allocate objects in the global heap. Given that a server can have cached objects (to accelerate reads), DRust carefully crafts an ownership-based cache coherence protocol upon the global heap abstraction to achieve both memory coherence and efficiency (§4.1.1).
@@
In a nutshell, our coherence protocol leverages the ownership semantics to eliminate the need for explicit cache invalidation. It allows multiple readers to fetch a copy of the object from its host server and cache it, but disallows any change to the global address and the value of the object. When a write access occurs, it must first borrow the ownership, at which point DRust moves the object in the global heap to a new address on the server issuing the write. The address change of the object automatically invalidates cache copies that use the stale address and triggers the subsequent readers to update the cache by fetching the object from its latest address.
@@
The second challenge is how to support transparency in programming. Rust’s standard libraries and programs were originally built for running on a single machine, and they cannot deal with distributed resources in a cluster. For example, a Rust program running on server A cannot spawn a thread on another server B, let alone synchronize threads between A and B. To enable a Rust program to run as is under DRust, we provide distributed threading utilities by restructuring critical elements of the Rust standard library, including threading, communication channels, and shared-state locks (§4.1.2). Our adapted libraries offer the same interfaces, making them compatible with single-machine Rust programs, but internally invoke our distributed scheduler, which determines where to run the thread and facilitates cross-server synchronization. We built them atop the ownership-based memory model, enabling the DRust runtime to safely pass references of objects between threads and automatically fetch the value from the global heap upon dereferencing.
@@
With our programming abstractions, a Rust application can start on a single server and gradually spawn its threads to other servers. Under the hood, DRust employs a runtime to manage distributed physical compute and memory resources for the application. The runtime runs as a process on each node in the cluster, and they work cooperatively for cross-server memory allocation and thread scheduling. The runtime prioritizes the current server for object allocation and thread creation, but it will schedule the resource allocation request to another server under memory pressure (§4.2.1). To make cluster-wise decisions such as deciding the target server for global memory allocation and thread creation, DRust has a global controller that is launched together with the application. The global controller communicates with DRust runtime on each node to collect resource usage information and applies adaptive policies to achieve load balance (§4.2.2).
@@
Results. We evaluated our system on four real-world applications in an eight-node cluster. Our evaluation demonstrated an average of 2.02$\times$ and 9.48$\times$ (up to 2.64$\times$ and 29.16$\times$) speedup compared with two state-of-the-art DSM systems GAM and Grappa, respectively. Furthermore, DRust incurred a mere 2.42% slowdown compared to the original Rust program on a single machine with sufficient resources. DRust is available at https://github.com/uclasystem/DRust.
@@ 4 Design
Figure 2: Design overview of DRust.
DRust is an efficient DSM framework atop the Rust programming language. As shown in Figure 2, it consists of Rust-based programming abstractions for DSM (§4.1) and a runtime (§4.2) that manages distributed physical resources.
DRust is compatible with standard Rust. Listing 2 illustrates how the accumulator (shown in Listing 1) runs on DRust distributively without requiring code rewriting. The program starts running on a single machine A and the DRust runtime gradually allocates its memory and spawns new threads on different machines. Specifically, Lines 10LABEL:lst:dist-acc:line:acc-end create Accumulator a and b where a.val and b are in the global heap. We use a global allocator to allocate objects in the global address space and hence these objects may be allocated on a different server. Line 15 synchronously adds b to a by fetching both values a.val and b to A’s local memory (if they are allocated somewhere else). Line LABEL:lst:dist-acc:line:remote-add spawns a new thread and ships the function closure to perform add asynchronously. This thread will be scheduled on a different server B if A’s compute power has been saturated. In this case, DRust performs shallow copying and only ships the pointers stored in a and b to B without actually moving objects in the global heap. The newly-created thread relies on the DRust runtime to detect data locations and fetch objects upon dereferencing.
@@
Listing 2: DRust seamlessly transforms an unmodified accumulator implemented in Rust into a distributed version.
4.1 DRust Programming Abstraction
DRust provides each thread with a local stack and abstracts distributed memory as a shared global heap. Each server allocates thread stacks and backs one partition of the global heap with its physical memory. DRust re-implemented core memory management constructs including Box, &, and &mut for transparent heap access. This approach hides the complex details of memory allocation/deallocation, moving objects, and coherence maintenance (§4.1.1). DRust supports distributed threading and synchronization by adapting Rust’s standard libraries atop the core language constructs (§4.1.2). Furthermore, DRust offers affinity annotations that allow developers to build more efficient applications by expressing data affinity semantics (§4.1.3).
@@ 4.1.1 Memory Management
Figure 3: The address space layout of DRust. The stack is private to each thread but they share an aligned address space to ease migration, while the heap is globally shared and partitioned across servers.
Next, we discuss how DRust (re)implements the memory-related language constructs in Rust to achieve memory safety and memory coherence.
Address Space. As shown in Figure 3, DRust maintains an identical address space layout on all servers. It exposes distributed memory as a coherent shared heap to applications. Embracing the idea of partitioned global address space (PGAS) [21], it partitions the heap space and assigns each server a unique address range. The stack, in contrast, is private to each thread. However, DRust aligns the stack space on each server and pads stacks to avoid overlapping. This streamlines thread migration between servers as it allows a thread to keep its private stack address unchanged when being moved.
Coherence Protocol in a Nutshell. For efficiency, DRust employs a call-by-reference model for newly created threads. Upon creation of a thread, the DRust runtime only passes references or Box pointers to objects to the newly created thread. Upon dereferencing, objects are fetched to the server where the thread is executed.
@@
Note that this is a general protocol that covers the case that the object is on the same server that issues the write—as long as the server moves the object into a different location in the global heap, no other servers can read the stale copies of the object. However, this is not efficient as each local write requires moving the object to a new address. To address this inefficiency, DRust employs a pointer-coloring technique, inspired by the designs of many concurrent garbage collectors [1, 52]. Discussed at the end of this subsection, this technique offers a more efficient solution for handling local writes.
Figure 4: DRust repurposes Rust pointers and references to contain a global heap address and an extension field for its coherence protocol.
Pointer Layout. In order to support this protocol, each pointer must remember not only the object’s global address, but also the address of the cached copy in a server’s local cache (to avoid redundant remote fetches). As such, we modify Rust’s pointer structure, as illustrated in Figure 4. DRust internally extends each Rust Box pointer and reference with an additional 64-bit field, which is used differently for read and write access. At a high level, the field records the address of the cached copy for faster read accesses; for write accesses, this field records the address of the object’s owner for post-write synchronization. Additionally, DRust reserves the highest 16 bits in the global address field as “color” bits. These bits record the version number of the pointer and play a crucial role in DRust’s efficient handling of local writes.
Next, we discuss how DRust reimplements Rust’s ownership operations to realize the distributed coherence protocol. For ease of presentation, this subsection focuses on a simplified version of the protocol. A complete coherence protocol and its proof of memory coherence are available in [53].
@@ 2 Function DerefMut($m$):
Mutable Borrow. Mutable borrow creates a mutable reference that holds exclusive access to the referenced object for writing. Algorithm 1 outlines the procedures for both dereferencing and dropping a mutable reference. When performing dereferencing, DRust first checks the object’s location (Line 1) and performs direct access if the object’s address belongs to the heap partition of the machine A that executes the access. Otherwise, DRust moves it to A’s heap partition (as opposed to caching it) (Line 1). The move, conducted in the following three steps, changes the object’s global address. DRust (1) copies the object into A’s heap at an address $p$, (2) updates the mutable reference with the address $p$, and (3) asynchronously requests the remote server that previously stored the object to deallocate the original object.
A challenge arises with its original owner Box, which now becomes a dangling pointer, pointing to an invalid memory location. Fortunately, the integrity of the system is maintained by the single-writer invariant (referenced as Invariant 3). This invariant ensures that while the mutable reference remains alive, no other entity, including the original owner, can access the data. To ensure correctness, when this new reference is dropped, DRust synchronously updates the original owner Box, redirecting it to the new address $p$ (Line 1). As a result, the original owner always possesses the latest view of the object. Additionally, all modifications made through this mutable reference are visible in all subsequent accesses, as they necessitate borrowing permission from the updated owner Box. The single-writer invariant also eliminates the possibility of simultaneous updates to the owner, ensuring that updating the owner is free from concurrency issues.
@@ 26             Atomic {
Immutable Borrow. Immutable borrowing allows concurrent reads to the same object from immutable references on the same or different servers. As detailed in Algorithm 2, DRust handles the dereferencing of immutable references by first checking the object’s location (Line 2). For remote objects, DRust creates a local copy in the per-node read-only “cache” and records its local address in the reference’s extension field (see Figure 4). This preserves the original global address of the object, ensuring that any new immutable reference—whether it is derived from the owner Box or from another immutable reference—can always access the original object from the global heap.
@@
As opposed to being a separate memory space, our “cache” provides a “virtual” aggregation of all local copies maintained on each server. These copies reside in the regular heap, managed by a per-node hashmap $H$. This hashmap maps each global address to a pair of its local address and the number of local immutable references to the local copy. To prevent redundant copies of an object on the same server, DRust checks the hashmap $H$ before creating a new local copy (Line 2). If a local copy is already present, DRust increments its reference count in $H$ and updates the extension field in the immutable reference to point to this copy (Lines 22). If no existing copy is found, a new one is created (Lines 22). Since the hashmap uses objects’ global addresses as keys, if an object has been modified by another server since its last read, its global address must have changed, making cache lookup fail even if a (stale) local copy exists.
@@
DRust actively updates the reference count of each local copy when an immutable reference is either dereferenced or dropped, as outlined in Lines 2 and 2. Utilizing these counts, the DRust runtime periodically scans the “cache” and lazily reclaims unreferenced copies (i.e., those with a zero reference count) under memory pressure (§4.2.1). This mechanism, in conjunction with the safe borrowing invariant (2), prevents the local cache from memory leaks or illegal accesses.
@@
Owner Access without Borrow. DRust treats a direct memory access via the owner Box as a pair of mutable/immutable borrow and return. Depending on the reference type, DRust uses the extension field of Box accordingly and executes the read/write dereferencing logic. A special case arises when a mutable owner is immutably borrowed and becomes immutable until all borrowed references return. In this case, the owner can only cache the object during the borrow and delay the move until the borrow finishes. This would not create any correctness issues because the owner cannot be used for write access during this period.
@@
Ownership Transfer. Similar to Rust, DRust does not move the actual value during the transfer and only copies the Box pointer. DRust additionally checks and resets the pointer’s extension field and frees the cached copy in the executing machine’s cache to avoid cache leakage.
@@
Memory Deallocation. Like Rust, DRust tracks the lifetime of an object via its owner. Given that ownership transfer is implemented by only evicting the cached copy of the object (without changing its global presence), the memory safety of DRust’s global heap is preserved by the singular owner invariant (1). In other words, DRust still guarantees that when an object’s owner goes out of scope, the object must be unreachable (and dead) and can be safely deallocated.
@@
Consistency Model. Our protocol, together with Rust’s ownership model, offers sequential consistency for cross-server memory accesses in safe Rust programs (i.e., following the original Rust, no guarantees can be provided when Rust Unsafe is used), which is a strong consistency order. Therefore, it allows any safe Rust program to preserve its memory consistency on DSM. Sequential consistency necessitates a coherent memory system, requiring not only the SWMR invariant but also the data-value invariant [57]. In simple terms, the data-value invariant requires that the latest write to a value is immediately visible to subsequent readers. As discussed earlier, DRust’s protocol moves an object upon a write and updates the owner immediately. Therefore, the latest value is globally visible after each mutable borrow finishes. Subsequent read accesses, either in the Owned state or the Shared state, are hence guaranteed to see the moved object and read its latest value.